Weight. The weight F of a body varies inversely with the square of the distance d between the body and the center of the Earth. If a man weighs 300 pounds 4,000 miles from the center of the Earth, how much will he weigh at a distance of 5,000 miles from the center of the Earth.
Answer by Boreal(15235)

F=k/d^2
300=k/(4*10^3)^2=k/1.6*10^7
k=4.8*10^9
F=4.8*10^9/(5*10^3)^2=4.8*10^9/2.5*10^7
=1.92*10^2
=192 pounds.